OpenStudy (beaubrun):

Density Questions? 1. What is the volume of a sample of helium that has a mass of 1.34 × 10^−3 g, and a density of 0.17847 g/L? Answer in units of mL 2. What is the density of a piece of metal that has a mass of 8.52 × 105 g and is 92 cm × 48.8 cm × 97.7 cm? Answer in units of g/cm^3 3. The density of gold is 19.3 g/cm^3. What is the volume of a sample of gold with mass 1.709 kg? Answer in units of cm3 4. If this sample of gold is a cube, how long is each edge? Answer in units of cm

OpenStudy (joannablackwelder):

Density = mass/volume

OpenStudy (anonymous):

1. (1.34 × 10^−3 g) / ( 0.17847 g/L) = 0.00751 L = 7.51 mL 2. (8.52 × 10^5 g) / (92 cm × 48.8 cm × 97.7 cm) = 1.94 g/cm^3 3. (1709 g) / (19.3 g/cm^3) = 88.5 cm^3 4. Take the cube root of the answer in 3: 4.46 cm